# On Touchard’s Identity: Generalizations and Related Results

### v1

@@ Abstract
Keywords: Catalan number, Touchard’s identity, harmonic number, Stirling number of the second kind, binomial transform.
@@ 1 Introduction
Identity (2) has further interesting immediate consequences. Evaluation at $x=2$ produces the self-inverse sequence binomial transform identity
@@
which was reported by Donaghey [3]. Setting x=1x=-1 and $x=3$, in turn, in (2), produces the following interesting identity:
@@
We will make frequent use of the following identity
@@
extending harmonic numbers to half-integer arguments and which is a consequence of the link between harmonic numbers, odd harmonic numbers and the digamma function.
@@ 2 A generalization of Touchard’s identity and related results
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}y^{r-s+2k}\left({1-y}\right)^{s-1}}=\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}y^{k+r-s}\left({1-y}\right)^{n-k+s-1}},$ (20)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+1}}}\,C_{k}=\frac{1}{{n+1}}\sum_{k=0}^{n}{2^{-k}C_{k+1}}$ (28)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k+1}}}\,C_{k}=\frac{{2^{2n}}}{{2n+1}}\binom{{2n}}{n}^{-1}\sum_{k=0}^{n}{2^{-3k}\binom{{2k}}{k}C_{k+1}}.$ (29)
@@ 4 Related results via binomial transform
Two sequences of complex numbers $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, are called a binomial transform pair if
@@
Use the binomial transform pair $\{a_{k}=1,\alpha_{k}=\delta_{k0}\}$ in (66). ∎
@@
where the binomial transform pair $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where
@@
we recognize the binomial transform pair $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where

### v2

@@ Abstract
Keywords: Catalan number, Touchard’s identity, Knuth’s old sum, harmonic number, Stirling number of the second kind, binomial transform.
@@ 1 Introduction
Identity (2) has further interesting immediate consequences. Evaluation at $x=2$ produces the following self-inverse sequence binomial-transform identity:
@@
which was reported by Donaghey [3]. Setting x=1x=-1, and $x=3$, in turn, in (2), produces the following interesting identity:
@@
We will make frequent use of the following identity:
@@
which extends harmonic numbers to half-integer arguments and which is a consequence of the link between harmonic numbers, odd harmonic numbers and the digamma function.
@@ 2 A generalization of Touchard’s identity and related results
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}x^{r-s+2k}\left({1-x}\right)^{s-1}}=\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}x^{k+r-s}\left({1-x}\right)^{n-k+s-1}},$ (20)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+1}}}\,C_{k}$ $=\frac{1}{{n+1}}\sum_{k=0}^{n}{2^{-k}C_{k+1}},$ (28)
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k+1}}}\,C_{k}$ $=\frac{{2^{2n}}}{{2n+1}}\binom{{2n}}{n}^{-1}\sum_{k=0}^{n}{2^{-3k}\binom{{2k}}{k}C_{k+1}},$ (29)
@@
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k-1}}\,C_{k}}=2^{2n}\binom{{2n}}{n}^{-1}\left({2\sum_{k=1}^{n}{2^{-3k}C_{k-1}C_{k+1}}-1}\right).$ (30)
@@
Identity (30) is derived by shifting the summation index on the right-hand side of (27), setting r=1/2r=-1/2 and using
$\binom{{k-1/2}}{{k+1}}=-\frac{{C_{k}}}{{2^{2k+1}}}.$ (34)
@@ 4 Related results via binomial transform
Two sequences of complex numbers $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, are called a binomial- transform pair if
@@
Use the binomial-transform pair $\{a_{k}=1,\alpha_{k}=\delta_{k0}\}$ in (68). Here $\delta_{ij}$ is Kronecker’s delta symbol. ∎
@@
where the binomial-transform pair $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, with
@@
we recognize the binomial-transform pair $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where
@@
5 Additional results
The methods described in this paper can also be used to derive generalizations of another well-known combinatorial identity, Knuth’s old sum or the Reed Dawson identity:
[…]
Identity (82) occurs at $x=0$ in the following polynomial identity [2, Equation (6)]:
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}\left({1-x}\right)^{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}x^{n-2k}}.$ (83)
Starting with the following variation on (83):
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}x^{k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}x^{2k}\left({1-x}\right)^{n-2k}},$ (84)
and proceeding as in the proof of Theorem 1 yields the following generalization of Knuth’s old sum.
Theorem 4.
If $r$ and $s$ are complex numbers that are not negative integers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}\binom{{k+r}}{s}^{-1}}=\frac{s}{{n+r}}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}\binom{{n+r-1}}{{2k+r-s}}^{-1}}.$ (85)
Equation (85) reduces to Knuth’s old sum (82) at $s=0$, while at $r=1=s$ we obtain the following binomial-transform identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}C_{k}}=2^{-n}\binom{{n}}{{\left\lfloor{n/2}\right\rfloor}}$, (86)
after using the fact that
$\sum_{r=0}^{m}{2^{-2r}\binom{{2r}}{r}}=\sum_{r=0}^{m}{\binom{{r-1/2}}{r}}=\binom{{m+1/2}}{m}=\left({2m+1}\right)2^{-2m}\binom{{2m}}{m}$. (87)
Remark 3.
Identity (86) was derived on the MathWorld website [5].
Choosing $r=0$ and $s=1/2$ in (85) and using
$\binom{k}{1/2}=\frac{2^{2k+1}}{\pi}\,\binom{2k}{k}^{-1}$ (88)
and
$\binom{{n-1}}{{2k-1/2}}=\frac{{2^{2n}}}{{n\pi}}\binom{{n}}{{2k}}\binom{{4k}}{{2k}}^{-1}\binom{{2\left({n-2k}\right)}}{{n-2k}}^{-1},$ (89)
gives our next result.
Proposition 11.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-3k}\binom{{2k}}{k}^{2}}=\frac{1}{{2^{2n}}}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{2^{-2k}\binom{{2k}}{k}\binom{{4k}}{{2k}}\binom{{2\left({n-2k}\right)}}{{n-2k}}}$. (90)
The following variation on Equation (83):
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}x^{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}\left(1-x\right)^{n-2k}},$ (91)
and the application of Lemma 3 leads to another generalization of Knuth’s old sum.
Theorem 5.
If $n$ and $r$ are non-negative integers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}\left({n-k}\right)^{r}}$ (92)
[…]
Next, using (91) and Lemma 3, we present a related result to Knuth’s old sum involving a binomial-transform pair.
Theorem 6.
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}a_{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}\alpha_{n-2k}}.$ (94)
Finally, by considering the following variation on identity (69):
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}C_{k+1}x^{k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}x^{2k}\left({1-x}\right)^{n-2k}},$ (95)
and working as in the proof of Theorem 1, we derive the next combinatorial identity.
Theorem 7.
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}C_{k+1}\binom{{k+r}}{s}^{-1}}=\frac{s}{{n+r}}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\binom{{n+r-1}}{{2k+r-s}}}^{-1}.$ (96)
In particular, setting $r=2$, $s=3/2$ and using the fact that
$\binom{k+2}{3/2}=\frac{16}{3\pi}\frac{2^{2k}}{C_{k+1}}$ (97)
and
$\binom{{n+1}}{{2k+1/2}}=\frac{{2^{2n+2}}}{\pi}\frac{{n+1}}{{\left({2n-4k+1}\right)\left({4k+1}\right)}}\binom{{n}}{{2k}}\binom{{4k}}{{2k}}^{-1}\binom{{2\left({n-2k}\right)}}{{n-2k}}^{-1},$ (98)
we obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-3k}C_{k+1}^{2}}$ (99)
$\qquad=\frac{1}{{\left({n+1}\right)\left({n+2}\right)2^{2n-1}}}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\left({2n-4k+1}\right)\left({4k+1}\right)2^{-2k}C_{k}\binom{{4k}}{{2k}}\binom{{2\left({n-2k}\right)}}{{n-2k}}}$. (100)